# Robust Preference for Dynamical Dark Energy in DESI BAO and SN Measurements

### v1

@@ II Dynamical Dark Energy Models
Focusing on flat models where the DE EoS $w(a)$ can be described by a continuous function of the scale factor, the first Friedman equation reads
@@
$+\rho_{\rm DE,0}\,a^{-3}\exp\left(-3\int_{a_{0}=1}^{a}\frac{w_{\rm DE}(a^{\prime})}{a^{\prime}}\,da^{\prime}\right)\bigg{]}.$ (1)
@@
When it comes to cosmological perturbations, diffeomorphism invariance of GR requires fixing a gauge. In the synchronous gauge, the line element reads Ma and Bertschinger (1995)
$ds^{2}=a^{2}(\tau)\left[-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\right],$ (2)
where $d\tau=dt/a(t)$ is the conformal time, $\delta_{ij}$ and $h_{ij}$ are the unperturbed and perturbed spatial part of the metric tensors. For a fluid component $i$, the equations governing the dimensionless density perturbations $\delta_{i}=\delta\rho_{i}/\rho_{i}$ and the divergence of the $i$-th fluid velocity $\theta_{i}=i\kappa^{j}v_{j}$ in the Fourier space are Ma and Bertschinger (1995):
$\delta^{\prime}_{i}$ $=$ $-(1+w_{i})\,\left(\theta_{i}+\frac{h^{\prime}}{2}\right)-3\mathcal{H}\left(\frac{\delta P_{i}}{\delta\rho_{i}}-w_{i}\right)\delta_{i}$ (3)
@@
$\theta^{\prime}_{i}$ $=$ $-\mathcal{H}\left(1-3\frac{\delta P_{i}}{\delta\rho_{i}}\right)\theta_{i}+\frac{\delta P_{i}/\delta\rho_{i}}{1+w_{i}}\,{\kappa}^{2}\,\delta_{i}-{\kappa}^{2}\sigma_{i},$ (4)
@@
II.1 Chevallier-Polarski-Linder parametrization
The model proposed by Chevallier, Polarski and Linder Chevallier and Polarski (2001); Linder (2003) (CPL hereafter) can be regarded as the baseline parameterization used in most analyses focusing on DDE, including this work. In this scenario, the DE EoS reads
$w_{\rm DE}(a)=w_{0}+w_{a}\times\left(1-a\right).$ (5)
@@ II.1 Chevallier-Polarski-Linder parametrization
II.2 Exponential parametrization
As a next step, we consider the exponential form for the DE EoS Dimakis et al. (2016); Pan et al. (2020b)
$w_{\rm DE}(a)=(w_{0}-w_{a})+w_{a}\times\exp\left(1-a\right).$ (6)
Up to the first order of the Taylor expansion, this description reduces to the CPL parameterization around $a\simeq a_{0}\equiv 1$. However, as $a$ moves far away from 1, the exponential form can introduce (small) deviations from the linear CPL regime without increasing the dimensionality of the parameter space Najafi et al. (2024).
II.3 Jassal-Bagla-Padmanabhan parametrization
The third parameterization studied in this work is the model proposed by Jassal-Bagla-Padmanabhan in Ref. Jassal et al. (2005) (JBP hereafter). In this case the DE EoS is
$w_{\rm DE}(a)=w_{0}+w_{a}\times a\left(1-a\right).$ (7)
@@ II.4 Logarithmic parametrization
$w_{\rm DE}(a)=w_{0}-w_{a}\times\ln a.$ (8)
To the best of our knowledge, this parameterization was originally introduced by G. Efstathiou in Ref. Efstathiou (1999) to capture the behaviour of a wide class of potential scalar field models of DE at low redshift $z\lesssim 4$. Here, with a fair amount of courage and following thorough stability tests, we extend this parameterization all the way up to $z\to\infty$. In principle, for some combinations of parameters, the logarithmic term can actually grow in absolute value and cause instabilities. [^7] To address this problem, in Ref. Yang et al. (2021b), some of us adopted flat priors $w_{a}\in[-3,0]$, thus removing positive values of $w_{a}$ a priori. In this work, we have performed additional stability tests, which revealed that numerical instabilities arise only when CMB data are considered on their own (we do not report these results for any of the models under study, as they are not informative). Without late-time data, DE parameters remain essentially unconstrained, and $w_{a}$ can acquire large positive values. Despite the logarithmic nature of the equation of state, for these values, the DE energy density does not remain negligible in the early Universe, triggering warnings and errors in the Boltzmann solver code CAMB Lewis et al. (2000); Howlett et al. (2012). In contrast, when late-time data are included, the DE parameters are significantly constrained in the region $w_{a}<0$, and no large positive values are allowed. As a result, in the allowed region of parameter space, the DE energy density remains negligible at early times and does not affect early Universe cosmology. However, given the current data constraints and the slow logarithmic growth, this is not the case. We find that the parameterization can be safely extended to high redshift because the DE contribution remains largely negligible compared to other components in the Universe’s energy budget.
II.5 Barboza-Alcaniz parametrization
The last (but as we shall see, not least) model involved in our analysis is the one proposed by Barboza and Alcaniz in Ref. Barboza and Alcaniz (2008) (referred to as BA hereafter). In this case, the dark energy equation of state (EoS) is characterized by the following functional form:
$w_{\rm DE}(a)=w_{0}+w_{a}\times\frac{1-a}{a^{2}+(1-a)^{2}}.$ (9)
@@ III Methods
III.1 Statistical Analyses
The cosmology resulting from all the five DDE models listed in Sec. II can be characterized by 7 free parameters: the physical baryon energy density $\Omega_{\mathrm{b}}h^{2}$, the physical cold dark matter energy density $\Omega_{\mathrm{c}}h^{2}$, the amplitude of the primordial scalar spectrum $A_{\mathrm{s}}$, its spectral index $n_{s}$, the optical depth to reionization $\tau$, the angular size of the sound horizon $\theta_{\rm{MC}}$, and the two free parameters describing the DE sector — i.e., the present-day value of the DE EoS $w_{0}$ and the parameter describing its dynamical evolution $w_{a}$. To compare the theoretical predictions against observations, we implement these models in five different modified versions of the publicly available cosmological code CAMB Lewis et al. (2000); Howlett et al. (2012) and explore the posterior distributions of the 7-dimensional parameter space by performing Markov Chain Monte Carlo (MCMC) analyses via the publicly available sampler Cobaya Lewis and Bridle (2002); Lewis (2013). We employ the fast dragging speed hierarchy implementation Neal (2005), which includes a convergence diagnostic following the Gelman-Rubin statistic Gelman and Rubin (1992). We adopt a threshold $R-1\lesssim 0.01$ for assessing chain convergence. In Tab. 1, we present the flat prior ranges on which the parameters are left to freely vary.
III.2 Datasets
@@
      Measurements of the power spectra of temperature and polarization anisotropies, $C_{\ell}^{TT}$, $C_{\ell}^{TE}$, and $C_{\ell}^{EE}$, at small scales ($\ell>30$), obtained by the Planck plik likelihood Aghanim et al. (2020b, d);
@@
      Measurements of the spectrum of temperature anisotropies, $C_{\ell}^{TT}$, at large scales ($2\leq\ell\leq 30$), obtained by the Planck Commander likelihood Aghanim et al. (2020b, d);
@@
      Measurements of the spectrum of E-mode polarization, $C_{\ell}^{EE}$, at large scales ($2\leq\ell\leq 30$), obtained by the Planck SimAll likelihood Aghanim et al. (2020b, d);
@@
      Reconstruction of the spectrum of the lensing potential, obtained by the Planck PR4 NPIPE data release Carron et al. (2022) used in combination with ACT-DR6 lensing likelihood Madhavacheril et al. (2024); Qu et al. (2024a). [^8] The NPIPE lensing map Carron et al. (2022) covers CMB angular scales in the range $100\leq\ell\leq 2048$ using the quadratic estimator and re-processing Planck time-ordered data with several improvements, including around 8% more data compared to the plik-lensing likelihood. Notice also that NPIPE and ACT-DR6 measurements explore distinct angular scales and overlap only across a portion of the sky. Additionally they have different noise levels and instrument-related systematics. Therefore they can be regarded as nearly independent lensing measurements.
  • DESI: baryon acoustic oscillations (BAO) measurements extracted by observations of galaxies & quasars Adame et al. (2024a), and Lyman-$\alpha$ Adame et al. (2024c) tracers from the first year of observations using the Dark Energy Spectroscopic Instrument (DESI). These include measurements of the transverse comoving distance, the Hubble horizon, and the angle-averaged distance as summarized in Tab. I of Ref. Adame et al. (2024b).
  • PantheonPlus: Distance moduli measurements of 1701 light curves of 1550 spectroscopically confirmed type Ia SN sourced from eighteen different surveys, gathered from the Pantheon-plus sample Scolnic et al. (2022); Brout et al. (2022).
  • DESY5: Distance moduli measurements of 1635 Type Ia SN covering the redshift range of $1.130.10<z<$ 1.13 that have been collected during the full five years of the Dark Energy Survey (DES) Supernova Program Abbott et al. (2024); Sánchez et al. (2024); Vincenzi et al. (2024), along with 194 low-redshift SN in the redshift range of $0.10.025<z<$ 0.1 which are in common with the Pantheon-plus sample Scolnic et al. (2022); Brout et al. (2022).
IV Results
@@
  • Table 2 and Figure 1 summarize the numerical constraints and parameter correlations for the baseline CPL case (5). The results for this case are detailed in Sec. IV.1.
  • Table 3 and Figure 2 present the results for the exponential parameterization (6), discussed in Sec. IV.2.
  • Table 4 and Figure 3 provide the results for the JBP EoS (7), discussed in Sec. IV.3.
  • Table 5 and Figure 4 summarize the results for the logarithmic parameterization (8), detailed in Sec. IV.4.
  • Table 6 and Figure 5 present the results for the BA parameterization (9), discussed in Sec. IV.5.
Conversely, readers interested in a comprehensive overview of the results, their interpretation, and implications can refer directly to Sec. V.
IV.1 Results for the CPL parameterization
@@
Figure 1: CPL parametrization (5) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
We recover all the results discussed in the recent DESI release paper Adame et al. (2024b). For Planck+DESI, the constraints favor a present-day quintessence EoS with $w_{0}=-0.44^{+0.34}_{-0.21}$ at 68% CL [^9] Hereafter, constraints will always be quoted at 68% CL unless otherwise specified., showing a notable shift away from $w_{0}=-1$. On the other hand, $w_{a}=-1.81^{+0.37}_{-1.1}$ provides hints of dynamical evolution towards the phantom regime.
The addition of PantheonPlus significantly refines the constraints on the parameter space, reducing the error bars on the DE parameters by up to a factor of 5. Although $w_{0}$ shifts towards $-1$, it remains strictly in the quintessence regime: $w_{0}=-0.876\pm 0.045$. Consistent with DESI 2024 findings Lodha et al. (2024), the mean value of $w_{a}$ increases compared to Planck+DESI, now reading $w_{a}=-0.51^{+0.20}_{-0.17}$. This boosts the evidence for a past phantom-like DDE component to above 95% CL, see also Fig. 1.
@@
$100\theta_{\mathrm{MC}}$ $1.04092\pm 0.00029$ $1.04097\pm 0.00030$ $1.04094\pm 0.00030$
$\tau$ $0.0530\pm 0.0073$ $0.0545\pm 0.0071$ $0.0538\pm 0.0074$
$n_{\mathrm{s}}$ $0.9655\pm 0.0038$ $0.9662\pm 0.0038$ $0.3161\pm 0.0066$
@@
$w_{0}$ $-0.44^{+0.34}_{-0.21}$ $-0.876\pm 0.045$ $-0.726\pm 0.069$
$w_{a}$ $-1.81^{+0.37}_{-1.1}$ $-0.51^{+0.20}_{-0.17}$ $-1.05^{+0.34}_{-0.28}$
$\Omega_{\mathrm{m}}$ $0.344^{+0.033}_{-0.027}$ $0.3088\pm 0.0067$ $0.3161\pm 0.0066$
$\sigma_{8}$ $0.791^{+0.021}_{-0.028}$ $0.8192\pm 0.0098$ $0.8130\pm 0.0091$
$S_{8}$ $0.846^{+0.016}_{-0.013}$ $0.8310\pm 0.0095$ $0.8345\pm 0.0097$
@@
Table 2: CPL Parametrization (5) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. III. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{CPL}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
IV.2 Results for the Exponential parametrization

### v2

@@ 2 Dynamical Dark Energy Models
Focusing on flat Friedmann-Lemaître-Robertson-Walker cosmology and models where the DE EoS $w(a)$ can be described by a continuous function of the scale factor, the first Friedman equation reads
@@
$+\rho_{\rm DE,0}\,a^{-3}\exp\left(-3\int_{a_{0}=1}^{a}\frac{w(a^{\prime})}{a^{\prime}}\,da^{\prime}\right)\bigg{]}.$ (2.1)
@@
When it comes to cosmological perturbations, diffeomorphism invariance of GR requires fixing a gauge. In the synchronous gauge, the line element reads [385]
$ds^{2}=a^{2}(\tau)\left[-d\tau^{2}+(\delta_{ij}+h_{ij})dx^{i}dx^{j}\right],$ (2.2)
where $d\tau=dt/a(t)$ is the conformal time, $\delta_{ij}$ and $h_{ij}$ are the unperturbed and perturbed spatial part of the metric tensors. For a fluid component $i$, the equations governing the dimensionless density perturbations $\delta_{i}=\delta\rho_{i}/\rho_{i}$ and the divergence of the $i$-th fluid velocity $\theta_{i}=i\kappa^{j}v_{j}$ in the Fourier space are [385]:
$\delta^{\prime}_{i}$ $=$ $-(1+w_{i})\,\left(\theta_{i}+\frac{h^{\prime}}{2}\right)-3\mathcal{H}\left(\frac{\delta P_{i}}{\delta\rho_{i}}-w_{i}\right)\delta_{i}$ (2.3)
@@
$\theta^{\prime}_{i}$ $=$ $-\mathcal{H}\left(1-3\frac{\delta P_{i}}{\delta\rho_{i}}\right)\theta_{i}+\frac{\delta P_{i}/\delta\rho_{i}}{1+w_{i}}\,{\kappa}^{2}\,\delta_{i}-{\kappa}^{2}\sigma_{i},$ (2.4)
@@
2.1 Chevallier-Polarski-Linder parametrization
The model proposed by Chevallier, Polarski and Linder [170, 172] (CPL hereafter) can be regarded as the baseline parameterization used in most analyses focusing on DDE, including this work. In this scenario, the DE EoS reads
$w(a)=w_{0}+w_{a}\times\left(1-a\right).$ (2.5)
@@ 2.1 Chevallier-Polarski-Linder parametrization
2.2 Exponential parametrization
As a next step, we consider the exponential form for the DE EoS [386, 387]
$w(a)=(w_{0}-w_{a})+w_{a}\times\exp\left(1-a\right).$ (2.6)
Up to the first order of the Taylor expansion, this description reduces to the CPL parameterization around $a\simeq a_{0}\equiv 1$. However, as $a$ moves far away from 1, the exponential form can introduce (small) deviations from the linear CPL regime without increasing the dimensionality of the parameter space [259].
2.3 Jassal-Bagla-Padmanabhan parametrization
The third parameterization studied in this work is the model proposed by Jassal-Bagla-Padmanabhan in Ref. [178] (JBP hereafter). In this case the DE EoS is
$w(a)=w_{0}+w_{a}\times a\left(1-a\right).$ (2.7)
@@ 2.4 Logarithmic parametrization
$w(a)=w_{0}-w_{a}\times\ln a.$ (2.8)
To the best of our knowledge, this parameterization was originally introduced by G. Efstathiou in Ref. [169] to capture the behaviour of a wide class of potential scalar field models of DE at low redshift $z\lesssim 4$. Here, with a fair amount of courage and following thorough stability tests, we extend this parameterization all the way up to $z\to\infty$. In principle, for some combinations of parameters, the logarithmic term can actually grow in absolute value and cause instabilities. [^10] To address this problem, in Ref. [313], some of us adopted flat priors $w_{a}\in[-3,0]$, thus removing positive values of $w_{a}$ a priori. In this work, we have performed additional stability tests, which revealed that numerical instabilities arise only when CMB data are considered on their own (we do not report these results for any of the models under study, as they are not informative). Without late-time data, DE parameters remain essentially unconstrained, and $w_{a}$ can acquire large positive values. Despite the logarithmic nature of the equation of state, for these values, the DE energy density does not remain negligible in the early Universe, triggering warnings and errors in the Boltzmann solver code CAMB [388, 389]. In contrast, when late-time data are included, the DE parameters are significantly constrained in the region $w_{a}<0$, and no large positive values are allowed. As a result, in the allowed region of parameter space, the DE energy density remains negligible at early times and does not affect early Universe cosmology. However, given the current data constraints and the slow logarithmic growth, this is not the case. We find that the parameterization can be safely extended to high redshift because the DE contribution remains largely negligible compared to other components in the Universe’s energy budget.
2.5 Barboza-Alcaniz parametrization
The last (but as we shall see, not least) model involved in our analysis is the one proposed by Barboza and Alcaniz in Ref. [181] (referred to as BA hereafter). In this case, the DE EoS is characterized by the following functional form:
$w(a)=w_{0}+w_{a}\times\frac{1-a}{a^{2}+(1-a)^{2}}.$ (2.9)
@@ 3 Methods
3.1 Statistical Analyses
The cosmology resulting from all the five DDE models listed in Sec. 2 can be characterized by 8 free parameters: the physical baryon energy density $\Omega_{\mathrm{b}}h^{2}$, the physical cold dark matter energy density $\Omega_{\mathrm{c}}h^{2}$, the amplitude of the primordial scalar spectrum $A_{\mathrm{s}}$, its spectral index $n_{s}$, the optical depth to reionization $\tau$, the angular size of the sound horizon $\theta_{\rm{MC}}$, and the two free parameters describing the DE sector — i.e., the present-day value of the DE EoS $w_{0}$ and the parameter describing its dynamical evolution $w_{a}$. To compare the theoretical predictions against observations, we implement these models in five different modified versions of the publicly available cosmological code CAMB [388, 389] and explore the posterior distributions of the 8-dimensional parameter space by performing Markov Chain Monte Carlo (MCMC) analyses via the publicly available sampler Cobaya [390, 391] that employs the fast dragging speed hierarchy implementation [392]. The convergence of the generated MCMC chains is assessed via the Gelman-Rubin parameter $R-1$ [393]. For all models and datasets, we require $R-1<0.01$ for the chains to be considered converged. In Tab. 1, we present the flat prior ranges on which the parameters are left to freely vary.
3.2 Datasets
@@
      Measurements of the power spectra of temperature and polarization anisotropies, $C_{\ell}^{TT}$, $C_{\ell}^{TE}$, and $C_{\ell}^{EE}$, at small scales ($\ell>30$), obtained by the Planck plik likelihood [40, 394];
@@
      Measurements of the spectrum of temperature anisotropies, $C_{\ell}^{TT}$, at large scales ($2\leq\ell\leq 30$), obtained by the Planck Commander likelihood [40, 394];
@@
      Measurements of the spectrum of E-mode polarization, $C_{\ell}^{EE}$, at large scales ($2\leq\ell\leq 30$), obtained by the Planck SimAll likelihood [40, 394];
@@
      Reconstruction of the spectrum of the lensing potential, obtained by the Planck PR4 NPIPE data release [395] used in combination with ACT-DR6 lensing likelihood [56, 269]. [^11] The NPIPE lensing map [395] covers CMB angular scales in the range $100\leq\ell\leq 2048$ using the quadratic estimator and re-processing Planck time-ordered data with several improvements, including around 8% more data compared to the plik-lensing likelihood. Notice also that NPIPE and ACT-DR6 measurements explore distinct angular scales, as ACT uses only CMB multipoles $3000600<\ell<$ 3000 and has only partial overlap with the 67% sky fraction used in the Planck analysis [269]. Additionally they have different noise levels and instrument-related systematics. Therefore they can be regarded as nearly independent lensing measurements.
  • DESI: Baryon acoustic oscillations (BAO) measurements extracted by observations of galaxies & quasars [58], and Lyman-$\alpha$ [396] tracers from the first year of observations using the Dark Energy Spectroscopic Instrument (DESI). These include measurements of the transverse comoving distance, the Hubble horizon, and the angle-averaged distance as summarized in Tab. I of Ref. [336].
  • PantheonPlus: Distance moduli measurements of 1701 light curves of 1550 spectroscopically confirmed type Ia SN sourced from eighteen different surveys, gathered from the Pantheon-plus sample [358, 55].
  • DESY5: Distance moduli measurements of 1635 Type Ia SN covering the redshift range of $1.130.10<z<$ 1.13 that have been collected during the full five years of the Dark Energy Survey (DES) Supernova Program [60, 61, 62], along with 194 low-redshift SN in the redshift range of $0.10.025<z<$ 0.1 which are in common with the Pantheon-plus sample [358, 55].
[…]
4 Results
@@
  • Table 2 and Figure 1 summarize the numerical constraints and parameter correlations for the baseline CPL case (2.5). The results for this case are detailed in Sec. 4.1.
  • Table 3 and Figure 2 present the results for the exponential parameterization (2.6), discussed in Sec. 4.2.
  • Table 4 and Figure 3 provide the results for the JBP EoS (2.7), discussed in Sec. 4.3.
  • Table 5 and Figure 4 summarize the results for the logarithmic parameterization (2.8), detailed in Sec. 4.4.
  • Table 6 and Figure 5 present the results for the BA parameterization (2.9), discussed in Sec. 4.5.
Conversely, readers interested in a comprehensive overview of the results, their interpretation, and implications can refer directly to Sec. 5. Instead, a comprehensive discussion of the behavior of the EoS inferred from the different datasets employed in the analysis across various models, as well as constraints on crucial quantities that aid in interpreting the results discussed in this section – such as the pivot redshift (i.e., the redshifts where the equation of state is better constrained by current data) and phantom crossing – is presented in Appendix A and summarized in Fig. 7 and Tab. 7. Interested readers can refer to this appendix for further details.
4.1 Results for the CPL parameterization
@@
$100\theta_{\mathrm{MC}}$ $1.04092\pm 0.00029$ $1.04098\pm 0.00030$ $1.04094\pm 0.00030$
$\tau$ $0.0530\pm 0.0073$ $0.0547\pm 0.0073$ $0.0538\pm 0.0074$
$n_{\mathrm{s}}$ $0.9655\pm 0.0038$ $0.9663\pm 0.0038$ $0.9659\pm 0.0038$
@@
$w_{0}$ $-0.44^{+0.34}_{-0.21}$ $-0.820\pm 0.064$ $-0.726\pm 0.069$
$w_{a}$ $-1.81^{+0.37}_{-1.1}$ $-0.77\pm 0.28$ $-1.05^{+0.34}_{-0.28}$
$\Omega_{\mathrm{m}}$ $0.344^{+0.033}_{-0.027}$ $0.3088\pm 0.0070$ $0.3161\pm 0.0066$
$\sigma_{8}$ $0.791^{+0.021}_{-0.028}$ $0.8186\pm 0.0094$ $0.8130\pm 0.0091$
$S_{8}$ $0.846^{+0.016}_{-0.013}$ $0.8304\pm 0.0097$ $0.8345\pm 0.0097$
@@
Table 2: CPL Parametrization (2.5) – 68% CL constraints on the free and derived cosmological parameters for 3 different data combinations detailed in Sec. 3. Negative values of $\Delta\chi^{2}=\chi^{2}_{\text{CPL}}-\chi^{2}_{\Lambda\text{CDM}}$ indicate an improvement in the fit to the data compared to $\Lambda$CDM.
Figure 1: CPL parametrization (2.5) – one-dimensional posterior distributions and two-dimensional marginalized contours for the main key parameters as obtained from the Planck+DESI, Planck+DESI+PantheonPlus, and Planck+DESI+DESY5 dataset combinations.
We recover all the results discussed in the recent DESI release paper [336]. For Planck+DESI, the constraints favor a present-day quintessence EoS with $w_{0}=-0.44^{+0.34}_{-0.21}$ at 68% CL, [^12] Hereafter, constraints will always be quoted at 68% CL unless otherwise specified. showing a notable shift away from $w_{0}=-1$. On the other hand, $w_{a}=-1.81^{+0.37}_{-1.1}$ provides hints of dynamical evolution towards the phantom regime.
[…]
4.2 Results for the Exponential parametrization